OpenStudy (anonymous):

In triangle QRS, angle Q =3x+20, angle R=2x+37 and angle S=4x+15. Determine the measure of each angle.

OpenStudy (anonymous):

The sum of all internal angles f triangle =180 degrees. so 3x+20+2x=37+4x+15=180 now solve this linear equation in 1 variable for x and substitute said value in all 3 angles to get their values